Pellets of mass 1.5 g are fired in parallel paths with speeds of 110 m/s through a hole 3.5 mm in diameter.How far from the hole must you be to detect a 1.0-cm-diameter spread in the beam of pellets?
Answer #1

m=1.5*10^-3kg
v=110m/s
d=3.5*10^-3 m

wavelength of the pellet

lamda=h/p=h/mv

lamda=6.626*10^-34 J s/1.510^-3kg*110m/s
lamda=4.015*10^-33 m

condition for diffraction maxima

Sin threta= 1.22lamda/D

for small angle
r=l sin theta= l tan theta=l1.22lamda/D

0.5*10^-2m= l*1.22*4.015*10^-33 m/(3.5*10^-3 m/2)
=1.53*10^-27 m